Maria is 6yrs older than John, John is twice as old as Michael? if all their ages are added together it could be 102 how old is maria
Answer by stanbon(75887)

Maria is 6yrs older than John,
John is twice as old as Michael?
If all their ages are added together it could be 102
how old is maria
-----------------------
Equations:
m = j + 6
j = 2k
m + j + k = 102
----
Substitute and solve for "j":
j+6 + j + j/2 = 102
---
(5/2)j = 96
j = (2/5)*96
j = 38.4
---
Maria = j+6 = 44.4 yrs old
----------------------
Cheers,
Stan H.
===============